OpenStudy (anonymous):

i need help with an algebra 2 question

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

Ok whats the question?

OpenStudy (anonymous):

\[\left( \sqrt[6]{x^5?} \right)^7\]

OpenStudy (anonymous):

i cant even see it lol

OpenStudy (anonymous):

how to transform \left( \sqrt[6]{x^5?} \right)^7 into an expression with a rational exponent

OpenStudy (anonymous):

Boooooblah do you think you could help me after?

OpenStudy (anonymous):